OpenStudy (anonymous):

In 1994, the average price of a new domestic car was $16,930. In 2002, the average price was $19,126. Based on a linear model, what would be the approximate predicted average price for 2008? $23,000 $21,300 $20,800 $18,600

OpenStudy (anonymous):

igot a am i right?

OpenStudy (anonymous):

i got 20773 but that's not one of the choices.

OpenStudy (anonymous):

so im right??

OpenStudy (anonymous):

i guess so. but B would be more reasonoable answer.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Sorry. by bad aggain. here i'll show you the steps. but me and you are wrong. ok the difference between 2002 and 1994 is 8 years. and 19126 and 16930 is $2196. so in 8 years the price has gone up only by 2196. we take 2196/8 = $274.5 so each year it goes up by 274.5 we need to find in 2008. right? 2008-1994=14. so that is in 14 years how much has it gone up by. take 274.5*14=3843 now 3843+16930=20773

OpenStudy (anonymous):

so it's C. not a or b.

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

The junior class sold 120 turkey dinner plates and 200 chicken dinner plates for a total of $2,150. The senior class sold 100 turkey plates and 300 chicken plates, raising $2,625. What was the cost of each turkey dinner plate? $6 $6.25 $7 $7.50 i got d